OpenStudy (anonymous):

write a quadratic equation using solutions -5 and -2 ?=0

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x + 5)(x + 2) = 0 }\) Now multiply x + 5 and x + 2.

OpenStudy (unklerhaukus):

\[(x+5)(x+2)=0\]\[x^2+7x+10=0\]

Parth (parthkohli):

I think you know how I got that...zero product property

OpenStudy (anonymous):

(x+5)(x+2) = x^2 + 7x + 10